xsq + (12)sq = 288
x=
Show math

Answer by Alan3354(69443)

Here's one that's similar.
===================================
(8)sq + (12)sq = xsq
--------
Use ^ (Shift 6) for exponents
8^2 + 12^2 = x^2
64 + 144 = x^2
x^2 = 208
x = +sqrt(208)
x = -sqrt(208)